If (x/2) = 3, then the value of 3x + 2 is (a) 20 (b) 11 (c) 13/2 (d) 8

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Option (a) is correct here
Consider the given equation (x/2) = 3
To find out the value of x.

x=3×2

x= 6
Then, the value of 3x + 2 will be,

= (3 × 6) + 2
= 18 + 2
= 20


Thanks